How will you explain the relationship between the angle of release to the 

 height of the projectile? How about angle of release and range before it 

 reaches the farthest distance?


1
Expert's answer
2021-06-03T09:18:21-0400

The best way to explain the relationship it to derive it. So, the height can be found as


"h=\frac{[v_y^2]}{2g}=\frac{[(v\sin\theta)^2]}{2g}=\frac{v^2\sin^2\theta}{2g}."

As we see,


"h(\theta)=\frac{v^2}{2g}\sin^2\theta."

Find the range:


"R=[v_x][t]=[v\cos\theta]\bigg[2\frac{[v_y]}{g}\bigg]=[v\cos\theta]\bigg[2\frac{[v\sin\theta]}{g}\bigg],\\\space\\
R(\theta)=\frac{2v^2\sin\theta\cos\theta}{g}=\frac{v^2\sin(2\theta)}{g}."


The height depends on the squared sine of angle of launch, the range depends on the sine of double angle of launch.